Grade 11Physical Chemistry

The frequency of the radiation is 6×10^14 Hertz calculate Wave length and wave number and energy in joule

Mohanraj
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear student
 
As we know
 
c = f * lambda
 
3 * 10^8 = 6 * 10^14 * lambda
 
Lambda = 0.5 * 10^-6 metres
 = 5000 A°
 
Regards
Arun (askIITians forum expert)